how do I solve (x^2-5x+5)^(x^2-9x+20)=1?
Answer by stanbon(75887)

how do I solve (x^2-5x+5)^(x^2-9x+20)=1?
-----------------
That exponent must be zero.
x^2-9x+20 = 0
(x-5)(x-4) = 0
x = 5 or x = 4
==================
Cheers,
Stan H.